2ί, 4 Ω 3Ω a 6 Ω 2.5A is Is3Ω b Determine/draw the Norton equivalent circuit to terminals ab
Answer #1

- 3 VA + 2 VA +LOVA 30 15VA = 30 2=مد To find w 22 Yur для 32 AAA 62 G TA 3° Apply kul to right loop. -vt (361) + 3ix & ix=toil+2 (12-13) + 412+ 3(12-13) = O => +61 +giz-513 = SO ii in loop (3 31 3+3 (13-12)=0 » -3i2 +6i3=0 ID Ti = 4A, 12=-4A, i3=5Sol ก: Given circuit zix 312 Дл- a 6023 2.5A We have to calculate Nortans equivalent circuit i.ee IN= Nortons equivalent cu